the planet mars orbits 2.279×10^8 km from the sun express the distance using the appropriate prefix and show your calculations
Answer #1

given distance = 2.279 x 10^8 km

1 km = 1000 m

2.279 x 10^8 km = 2.279 x 10^8 x 10^3 m

                           = 2.279 x 10^11 m

1 giga meter = 10^9 meters

2.279 x 10^11 m = 227.9 x 10^9 m

                            = 227.9 Gm